# Edge Universality for Inhomogeneous Random Matrices

### v1

@@ 7.1.5 Weighted $d$-regular graphs
In particular, equip each edge of $G_{N}$ with an independent Rademacher sign $\sigma_{ij}\in\{\pm 1\}$, and denote the signed adjacency by $\sigma\circ G_{N}$. Equivalently, this is the “new” part of the adjacency of the random 2-lift $\widetilde{G}_{N}$. By [BL06, Lemma 3.1],
@@
On the other hand, when $G_{N}$ is drawn uniformly at random, [He24] shows for $d\gg N^{2/3+\tau}$,
@@
We proceed by induction on the number of 2-lift steps $k$. For $k=1$, the claim follows directly from Corollary 7.7 and Eq (7.38).
@@ 7.1.6 Dimension-free deviation inequality
In [BvH24a], both small and large deviation inequalities were established for inhomogeneous Gaussian matrices under minimal sparsity assumptions. As a direct application of Corollary 4.7, in the undeformed case we recover Tracy-Widom fluctuations in the small-deviation regime, provided that assumptions (B1) and (B2) hold.
@@
This bound may be sharpened by a more detailed analysis of Theorem 4.4. Notably, in contrast to [BvH24a], there is no dimension factor $N$ in the tail probability bound. Such tail decay could extend to $\theta$-sub-Gaussian case by carefully analyzing the bound in Lemma 5.6.
@@ 7.1.7 GOE-GUE transition for IRM
We can also extend an interpolating ensemble with unequal real and imaginary parts, which has been studied by Feldheim and Sodin [FS10, Section V], to the inhomogeneous setting. Indeed, replace the GOE/GUE matrix $W$ in our definition 1.1 $H=\Sigma_{N}\circ W$ with the general Gaussian ensemble,
@@
we immediately we obtain GOE for T $\alpha=0$, $\alpha=1$ for GUE, and an anti-Symmetric GOE for $\alpha=\infty$( see [Meh04, Chapter 13] for the third ensemble).
@@
Let $H_{N}$ be an IRM matrix defined in Definition 1.1 but with real/complex Gaussian entries replaced by (7.45), the following limit for the largest eigenvalue $\lambda_{1}$ hold.
  1. 1.
@@
    $N^{\frac{2}{3}}(\lambda_{1}-2)\overset{D}{\rightarrow}\mathrm{TW}_{1};$ (7.47)
  2. 2.
@@
The argument follows essentially the same approach as in [FS10, Section V], with the sole modification being the adaptation to our inhomogeneous setting. ∎
@@ 7.1.8 IRM with quaternion entries
Our methods naturally extend to matrix models with quaternion entries ($\beta=4$). In particular, the edge eigenvalue analysis developed by Feldheim and Sodin for quaternion Wigner and Wishart ensembles [FS10, Section V] can be directly adapted to the inhomogeneous edge setting with only minor modifications.
@@ 7.2 Further questions
    Edge universality beyond Assumption (B1). Assumption (B1) is both essential and technical. Violating this assumption may lead to non-universal edge corrections in the limiting spectral distribution. Nevertheless, after accounting for such corrections, edge universality may still hold; see, for instance, [LZ24b] in the band matrix setting.
@@
    Moment condition and edge corrections. In Theorem 1.3, we require the moment condition $\theta t_{N}\ll N^{1/3}$ in order to establish edge universality. Violating this condition may also lead to non-universal corrections at the spectral edge. However, as in the previous case, edge universality may be still true after such corrections. This phenomenon has been observed in the setting of sparse random matrices; see e.g. [HY22]. We also refer the reader to [ABL22] for edge statistics results in the heavy-tailed setting.
@@
    Variance profile without the doubly stochastic condition. Assumption (A2) imposes a doubly stochastic structure on the variance profile, which guarantees that the limiting empirical spectral distribution is the semicircle law. This restriction, however, is somewhat artificial and may be only technical. Removing Assumption (A2) is a very interesting problem; nonetheless, local universality of spectral statistics may still hold without it. For related results in matrix models that do not possess a doubly stochastic structure but require comparable variance profile, see e.g. [AEK17] for bulk universality, [AEKS20] for edge universality and [DGH24] for a large deviation principle.
@@
    Delocalization of edge eigenvectors. This paper focuses solely on eigenvalue statistics. A natural further question is whether the eigenvectors associated with the edge eigenvalues are delocalized under Assumptions (B1) and (B2). We refer to [YY25b] for recent progress on edge eigenvector delocalization for band matrix models.
@@
    Inhomogeneous non-Hermitian random matrices. In Theorem 6.7, we establish the universality of the largest singular values for inhomogeneous non-Hermitian matrices. However, the precise conditions under which eigenvalue universality holds for inhomogeneous non-Hermitian matrices remain unclear; we refer to [CES21] for edge universality of non-Hermitian random matrices with i.i.d. entries and to [LZ24a] for phase transition of eigenvalues of deformed Ginibre ensembles.
@@ Acknowledgments
We thank Jiaqi Fan and Fan Yang for valuable discussions and for kindly explaining their work [SYY25, FSYY25] to us. We also thank Ruohan Geng for valuable conversations, particularly on the diagram expansion of the deformed Tracy-Widom law. We appreciate Jiaoyang Huang for valuable discussions on random lifts of $d$-regular graphs. We are thankful to Ramon van Handel for valuable discussions on the fluctuation scale at the spectral edge during Guangyi’s visit to Princeton. We are grateful to Yuanyuan Xu for valuable discussions on the challenge of edge statistics for non-mean-field random matrices. This work was supported by the National Natural Science Foundation of China under Grants #12371157 and #12090012.
@@ Appendix A Chebyshev expansion in deformed Hermitian case
Using [EK11a, Lemma 6.1] (note the are slight different in notation $\underline{\Phi_{3}V_{\ell_{0}-2}}$),
@@ Appendix B Chebyshev expansion in the Wishart Case
The proof is the same as that in [EK11a, Lemma 6.1]. ∎
@@
Proceeding inductively and following the argument in [EK11a, Proposition 6.2], we can obtain the claimed path expansion for $\mathcal{Q}_{n}(X)$. ∎
@@ Appendix D Upper bounds for transition probabilities
$\widehat{p}_{n}(k)$ $=\frac{1}{M}\sum_{x\in\Lambda}\sum_{n\in\mathbb{Z}^{d}}f\!\left(\frac{x+nW}{W}\right)e^{-\frac{2\pi ik\cdot x}{L}}=\frac{1}{M}\sum_{x\in\mathbb{Z}^{d}}f\!\left(\frac{x}{W}\right)e^{-\frac{2\pi ik\cdot x}{L}}$
@@
where in the last two equalities we have used Poisson summation and dropped $|n|>0$ terms via (7.14). In the Poisson summation step, both series converge absolutely by (7.14), hence uniformly. Furthermore, they define bounded continuous functions. By Plancherel identity they agree in the sense of $L^{2}$ distance, and since two continuous functions that are equal a.e. must coincide everywhere on $\mathbb{R}^{d}$, the identities hold pointwise.
@@
Next, we estimate the main sum. We have
@@
For $I_{2}$, Assumption 7.2(c) gives
$I_{2}\;\leq\;\frac{1}{L^{d}}\Bigl{(}\tfrac{100C_{K}L}{W}\Bigr{)}^{d}\rho^{n}\;\leq\;C^{\prime}_{K}W^{-d}\rho^{n}\;\leq\;C^{\prime\prime}_{K}W^{-d}n^{-\tfrac{d}{\alpha}}.$ (D.8)
For $I_{3}$, using Assumption 7.2(d),
@@
Combining these bounds with (D.6), we conclude the proof. ∎

### v2

@@ 7.1.6 Weighted $d$-regular graphs
In particular, equip each edge of $G_{N}$ with an independent Rademacher sign $\sigma_{ij}\in\{\pm 1\}$, and denote the signed adjacency by $\sigma\circ G_{N}$. Equivalently, this is the “new” part of the adjacency of the random 2-lift $\widetilde{G}_{N}$. By [MR2279667, Lemma 3.1],
@@
On the other hand, when $G_{N}$ is drawn uniformly at random, [MR4777082] shows for $d\gg N^{2/3+\tau}$,
@@
We proceed by induction on the number of 2-lift steps $k$. For $k=1$, the claim follows directly from Corollary 7.7 and (7.38).
@@ 7.1.7 Dimension-free deviation inequality
In [brailovskaya2024extremal], both small and large deviation inequalities were established for inhomogeneous Gaussian matrices under minimal sparsity assumptions. As a direct application of Corollary 4.7, in the undeformed case we recover Tracy-Widom fluctuations in the small-deviation regime, provided that assumptions (B1) and (B2) hold.
@@
This bound may be sharpened by a more detailed analysis of Theorem 4.4. Notably, in contrast to [brailovskaya2024extremal], there is no dimension factor $N$ in the tail probability bound. Such tail decay could extend to $\theta$-sub-Gaussian case by carefully analyzing the bound in Lemma 5.6.
@@ 7.1.8 GOE-GUE transition for IRM
We can also extend an interpolating ensemble with unequal real and imaginary parts, which has been studied by Feldheim and Sodin [feldheim2010universality, Section V], to the inhomogeneous setting. Indeed, replace the GOE/GUE matrix $W$ in our definition 1.1 $H=\Sigma_{N}\circ W$ with the general Gaussian ensemble,
@@
we immediately obtain GOE for T $\alpha=0$, $\alpha=1$ for GUE, and an anti-Symmetric GOE for $\alpha=\infty$( see [Mehta2004, Chapter 13] for the third ensemble).
@@
Let $H_{N}$ be an IRM matrix defined in Definition 1.1 but with real/complex Gaussian entries replaced by (7.45), the following limits for the largest eigenvalue $\lambda_{1}$ hold.
  1. (1)
@@
    $N^{\frac{2}{3}}(\lambda_{1}-2)\overset{D}{\rightarrow}\mathrm{TW}_{1}.$ (7.47)
  2. (2)
@@
The argument follows essentially the same approach as in [feldheim2010universality, Section V], with the sole modification being the adaptation to our inhomogeneous setting. ∎
@@ 7.1.9 IRM with quaternion entries
Our methods naturally extend to matrix models with quaternion entries ($\beta=4$). In particular, the edge eigenvalue analysis developed by Feldheim and Sodin for quaternion Wigner and Wishart ensembles [feldheim2010universality, Section V] can be directly adapted to the inhomogeneous edge setting with only minor modifications.
@@ 7.2 Further questions
    Edge universality beyond Assumption (B1). Assumption (B1) is both essential and technical. Its violation can introduce non-universal corrections at the spectral edge when analyzing Tracy-Widom fluctuations. Nevertheless, the core edge universality may remain intact, provided these corrections are incorporated; see [liu2023edge] for an example in the context of band matrices.
@@
    Moment condition and edge corrections. In Theorem 1.3, we require the moment condition $\theta t_{N}\ll N^{1/3}$ in order to establish edge universality. Violating this condition may also lead to non-universal corrections at the spectral edge. However, as in the previous case, edge universality may be still true after such corrections. This phenomenon has been observed in the setting of sparse random matrices; see e.g. [huang2022edge]. We also refer the reader to [aggarwal2022mobility] for edge statistics results in the heavy-tailed setting.
@@
    Variance profile without the doubly stochastic condition. Assumption (A2) imposes a doubly stochastic structure on the variance profile, which guarantees that the limiting empirical spectral distribution is the semicircle law. This restriction, however, is somewhat artificial and may be only technical. Removing Assumption (A2) is a very interesting problem; nonetheless, local universality of spectral statistics may still hold without it. For related results in matrix models that do not possess a doubly stochastic structure but require comparable variance profile, see e.g. [MR3719056] for bulk universality, [MR4089499] for edge universality and [ducatez2024large] for a large deviation principle.
@@
    Delocalization of edge eigenvectors. This paper focuses solely on eigenvalue statistics. A natural further question is whether the eigenvectors associated with the edge eigenvalues are delocalized under Assumptions (B1) and (B2). We refer to [yang2025delocalization2] for recent progress on edge eigenvector delocalization for band matrix models.
@@
    Inhomogeneous non-Hermitian random matrices. In Theorem 6.7, we establish the universality of the largest singular values for inhomogeneous non-Hermitian matrices. However, the precise conditions under which eigenvalue universality holds for inhomogeneous non-Hermitian matrices remain unclear; we refer to [MR4221653] for edge universality of non-Hermitian random matrices with i.i.d. entries and to [liu2024] for phase transition of eigenvalues of deformed Ginibre ensembles.
@@ Acknowledgments
We thank Jiaqi Fan and Fan Yang for valuable discussions and for kindly explaining their work [stone2025random, fan2025localization] to us. We also thank Ruohan Geng for valuable conversations, particularly on the diagram expansion of the deformed Tracy-Widom law. We appreciate Jiaoyang Huang for valuable discussions on random lifts of $d$-regular graphs. We are thankful to Ramon van Handel for valuable discussions on the fluctuation scale at the spectral edge during Guangyi’s visit to Princeton. We are grateful to Yuanyuan Xu for valuable discussions on the challenge of edge statistics for non-mean-field random matrices. This work was supported by the National Natural Science Foundation of China under Grants #12371157 and #12090012.
@@ Appendix A Chebyshev expansion in deformed Hermitian case
Using [EK11Quantum, Lemma 6.1] (note the are slight different in notation $\underline{\Phi_{3}V_{\ell_{0}-2}}$),
@@ Appendix B Chebyshev expansion in the Wishart Case
The proof is the same as that in [EK11Quantum, Lemma 6.1]. ∎
@@
Proceeding inductively and following the argument in [EK11Quantum, Proposition 6.2], we can obtain the claimed path expansion for $\mathcal{Q}_{n}(X)$. ∎
@@ Appendix D Upper bounds for transition probabilities
$\widehat{p}(k)$ $=\frac{1}{M}\sum_{x\in\Lambda}\sum_{n\in\mathbb{Z}^{d}}f\!\left(\frac{x+nW}{W}\right)e^{-\frac{2\pi ik\cdot x}{L}}=\frac{1}{M}\sum_{x\in\mathbb{Z}^{d}}f\!\left(\frac{x}{W}\right)e^{-\frac{2\pi ik\cdot x}{L}}$
@@
where in the last two equalities we have used Poisson summation and dropped $|n|>0$ terms via (7.14). In the Poisson summation step, both series converge absolutely by (7.14), hence uniformly. Furthermore, the same two series define bounded continuous functions. By Plancherel identity they agree in the sense of $L^{2}$ distance, and since two continuous functions that are equal a.e. must coincide everywhere on $\mathbb{R}^{d}$, the identities hold pointwise.
@@
Next, we proceed to estimate the main sum:
@@
For $I_{2}$, Assumption 7.2(3) gives
$I_{2}\;\leq\;\frac{1}{L^{d}}\Bigl(\tfrac{100C_{K}L}{W}\Bigr)^{d}\rho^{n}\;\leq\;C^{\prime}_{K}W^{-d}\rho^{n}\;\leq\;C^{\prime\prime}_{K}W^{-d}n^{-\tfrac{d}{\alpha}},$ (D.8)
while for $I_{3}$, Assumption 7.2(4) leads to
@@
Finally, combining these bounds with (D.6), we conclude the proof. ∎